OpenStudy (anonymous):

If x^4 = y^16, then y =

OpenStudy (anonymous):

you dont have any choices

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

i mean like a b c or d

OpenStudy (anonymous):

i have the answer i just need to know how and why of the process in obtaining the answer

OpenStudy (asnaseer):

it looks like you first need to learn about the various laws of exponents. I suggest you look here first: http://www.mathsisfun.com/exponent.html and then come back to this question. If you are still stuck after this then ping me.

OpenStudy (anonymous):

yea still stuck....

OpenStudy (asnaseer):

which parts do you not understand?

OpenStudy (anonymous):

when you asked what is the 4th root of x^4

OpenStudy (asnaseer):

if you are given \(x^a\) then the a'th root of this is given by:\[\left(x^a\right)^{\frac{1}{a}}=x^{a\times\frac{1}{a}}=x^1=x\]

OpenStudy (anonymous):

yea ... i've seen that before but im not gettin the logic of that.... why do you do that?

OpenStudy (asnaseer):

so the square root of \(x^2\) is given by:\[\left(x^2\right)^\frac{1}{2}=x^{2\times\frac{1}{2}}=x^1=x\]

OpenStudy (asnaseer):

square root can be written in two forms:\[\sqrt{x}\text{ or }x^{\frac{1}{2}}\]

OpenStudy (anonymous):

yea i know that part

OpenStudy (asnaseer):

what logic do you not get?

OpenStudy (anonymous):

so okay if theres 4 on the left side and 4^16 on the right how is it multiplied by 1/16 on both sides and why?

OpenStudy (asnaseer):

let me try to explain in another way...

OpenStudy (anonymous):

sorry i meant 4(1/16) and 16(1/16)

OpenStudy (asnaseer):

I assume you that:\[x^2=x\times x\]and\[x^3=x\times x\times x\]etc...

OpenStudy (anonymous):

yea i know that of course lol

OpenStudy (asnaseer):

good :) now square root is called the "inverse" of squaring and cube root is called the "inverse" of cubing. so, if you have \(x^2\), you can get back to \(x\) by taking its square root, i.e.:\[\sqrt{x^2}=(x^2)^\frac{1}{2}=x\]

OpenStudy (asnaseer):

similarly, if you have \(x^3\) you can get back yo \(x\) by taking its cube root, i.e.:\[\sqrt[3]{x^3}=(x^3)^\frac{1}{3}=x\]

OpenStudy (anonymous):

oh .... hollup damn i think u just reminded me of something so a \[\sqrt{x} \] is equal to \[x ^{1/2}\]

OpenStudy (asnaseer):

and, in general, if you have the a'th power of \(x\), then you can get back to \(x\) by taking its a'th root, i.e.:\[\sqrt[a]{x^a}=(x^a)^\frac{1}{a}=x\]

OpenStudy (asnaseer):

yes - what you just said is correct

OpenStudy (anonymous):

so even if there isn't a x^2 and just an x inside the square root it'll still be a x^1/2?

OpenStudy (asnaseer):

yes:\[\sqrt{x}=x^{\frac{1}{2}}\]

OpenStudy (anonymous):

oh okay lol

OpenStudy (asnaseer):

now, in your question you are given \(y^{16}\) and are asked to find \(y\)

OpenStudy (asnaseer):

so you effectively need to take the 16'th root of both sides of the equation you are given

OpenStudy (anonymous):

so it'll be 1/16 on both sides right?

OpenStudy (asnaseer):

perfect! :)

OpenStudy (anonymous):

is there another way to do this problem ? you don't make it (4)^2?

OpenStudy (asnaseer):

I am not sure what you are asking by saying (4)^2 ?

OpenStudy (anonymous):

making the 16 on the right side a (4)^2

OpenStudy (asnaseer):

you have:\[x^4=y^{16}\]now we take the 16'th root of both sides to get:\[(x^4)^{\frac{1}{16}}=(y^{16})^{\frac{1}{16}}\]

OpenStudy (anonymous):

but we're trying to yet y by itself so that's why we're multiplying it by 1/16 tho right?

OpenStudy (anonymous):

get not yet*

OpenStudy (asnaseer):

yes - except we are not "multiplying" by 1/16 - we are taking the 16'th root

OpenStudy (asnaseer):

the indexes on \(x\) and \(y\) are being multiplied

OpenStudy (asnaseer):

or exponents

OpenStudy (anonymous):

yea okay i think i get it now

OpenStudy (anonymous):

thanks !!

OpenStudy (asnaseer):

yw :)